Grade 10Wave Motion

Two transparent slabs having equal thickness but different refractive indices μ1 and μ2 are pasted side by side to form a composite slab. This slab is placed just after the double slit in a Young’s experiment so that the light from one slit in a young’s experiment so that the light from the other slit goes through the other material. What should be the minimum thickness of the slab so that there is a minimum at the point P0 which is equidistant from the slits?

Hrishant Goswami
12 Years agoGrade 10

1 Answer

Jitender Pal
12 Years ago
Sol. The change in path difference due to the two slabs is (μ base 1 – μ base 2)t (as in problem no. 16). For having a minimum at P base 0, the path difference should change by λ/2. So, ⇒ λ/2 = (μ base 1 – μ base 2)t ⇒ t = λ/2(μ base 1 – μ base 2)